1÷x-1÷x-2=3,x is not equal to 0,2

Grade:10

2 Answers

yaswanth
9 years ago
1/x-1/x-2=3
x-2-x/x(x-2)=2
-2/x2-2x=3
3x2-6x+2=0
x=1+1/31/2 ,1-1/31/2
kartheeek
9 years ago
1/x-1/x-2=2/x*2-x=3
3x^2-6x+2=0
x=1+1/3^1/2 and 1-1/1^1/2